My question is as follows:
QUESTION) Find the slope-intercept form of the equation of a line that passes through the given points. Given points: (7,3), (-8,3)
What I detemined so far is that I think I need to use the Slope Intercept form y=mx+b. I know that Y intercept is 3, but am stuck after this point.
Thanks for your time and help!
Armon

Answer by stanbon(75887)

Find the slope-intercept form of the equation of a line that passes through the given points. Given points: (7,3), (-8,3)
-----------------
Since the y-values are the same the equation is y=3.
----------------
If you would go thru the normal procedure of finding the equation based
on the two points you would get the following result:
slope = (3-3)/(-8-7) = 0/-15 = 0
-----------
since the form is y=mx+band since you have x,y, and m you can solve for "b".
3 = 0*7 + b
b = 3
----------------
Putting m=0 and b=3 into the form you get:
y = 0x+3
OR
y = 3
===================
Cheers,
stan H.